OpenStudy (anonymous):

Find the points at which the following curves intersect (correct to 2 decimal places) y=x^2+3x-10 x^2+y^2-2x-6y-6=0

OpenStudy (anonymous):

I'd try it by substituting the first equation into the second and solving for x.

OpenStudy (anonymous):

It's possible for a parabola and an ellipse to intersect in, at most, 4 places, so be sure you can handle solving a 4th-degree polynomial.

OpenStudy (anonymous):

4th -degree polynomial?

OpenStudy (anonymous):

ive subbed the first equation in the second, then i get stuck there

OpenStudy (anonymous):

Yes, it's possible that after doing the substitution that you'll get an equation with x^4

OpenStudy (anonymous):

Did you simplify as much as possible?

OpenStudy (anonymous):

yeah i think

OpenStudy (anonymous):

on the picture there is curve and point

OpenStudy (anonymous):

i got \[5x ^{2}+y ^{2}-20x-54=0\]

OpenStudy (anonymous):

From the graph ajay posted, you can see that the 4th-degree equation will only have 2 real solutions. Are you familiar with the techniques of using the rational root theorem and synthetic division?

OpenStudy (anonymous):

no, not really

OpenStudy (anonymous):

When you do the substitution, you should get an equation with only one variable.

OpenStudy (anonymous):

so i can sub y into y^2?

OpenStudy (anonymous):

It is impossible to solve a single equation that has 2 variables in it. Start here, and simplify: \[x^2+(x^2+3x+10)^2-2x-6(x^2+3x+10)-6\] This is merely equation one substituted into equation two.

OpenStudy (anonymous):

Yes, the first equation is solved for y, so you can use that expression for y as a substitute in the second equation.

OpenStudy (anonymous):

okay ill try the simplifying part

OpenStudy (anonymous):

how do you do x^2* 3x

OpenStudy (anonymous):

@CliffSedge ?

OpenStudy (anonymous):

The rest is going to be very difficult if you don't know how to multiply monomials. \[x^2*3x=3x^3\]

OpenStudy (anonymous):

I'm going to double check my simplifying, I was expecting some negative terms..

OpenStudy (asnaseer):

@SirMathy what methods have you been taught to solve this? e.g. maybe you are being asked to plot these graphs and see where they intersect, or maybe you have been asked to use numerical analysis (like the Newton-Raphson Method)?

OpenStudy (anonymous):

From that analysis, it looks like the two real solutions are irrational - that can be difficult to figure out by hand..

OpenStudy (anonymous):

my math teacher said to use substitution

OpenStudy (asnaseer):

in that case, when you substitute the first equation into the second you end up with:\[x^4+6x^3-16x^2-80x+154=0\]have you been taught how to solve such equations?

OpenStudy (anonymous):

I asked, and SirMathy says he doesn't know the rational root theorem or synthetic division..

OpenStudy (asnaseer):

then we are missing something here. are you sure you have the right question?

OpenStudy (anonymous):

Yeah, look dont worry, ill ask my math teacher about this question tommorrow Thankyou for your help anyway

OpenStudy (asnaseer):

ok - yw